Journalist Report – November 21st

The white structures of the campus are overlaid on a primordial landscape of pink, grey, orange and white. The Hab accommodates 6 crew in its upper level, but one of the
reasons it feels psychologically comfortable is the views out of the 5 upper level windows. These help us keep in touch with the daily cycle. We don’t worry about the neighbors looking in.

Above the top stair landing facing south is one of two large circular windows. This
looks directly down across a flat parking area in front of the Hab enclosed by mounds
and hillocks to the south and east, where the rovers and other vehicles park and can be
monitored. The RAM can be seen off to the right. This window is located above the main
entry airlock and stairs, though these are not visible from here. Looming above the
adjacent mounds on the distant horizon is Mount Henry. This is the highest landform
visible from the Hab, and has a blueish tinge, now capped by snow.
The second large window in the upper level living space faces east, and is located over
the couch. It overlooks the Greenhab and Science Dome, and also the approach to the
MDRS from the public road heading northwards into the wider adjacent public lands. This
allows us to monitor our EVA’s departing and returning to the campus, and also any
other traffic that may be in the area. A hill on the far side of the road has a
distinctive square rock at the summit and is known as ‘Zubrin’s Head’. The land in that
direction is generally flat, though accommodates large eroded, rounded mounds of
varying sizes and colors. Through this window we enjoy beautiful sunrises.
Two small square windows are located above the kitchen bench. One facing north-east provides a view of the tunnel to the Science Dome and to the Musk Observatory. The separate forms of the second observatory and solar array are also visible. The flat
plain continues north and west, with large hills on the horizon in the far distance,
and medium sized hills with distinctive colored sedimentary striping and curved
silhouettes in the middle distance to the north. We can see the rovers driving off the
north through this window too, and pick up their radios as they return. It was through
this window we saw the flashes of Andrew’s heliopgraph during our tests the first week of the expedition. The small square window above the sink catches a fuller view of the striped hills to the north, with the adjacent section of plain in the foreground.
My stateroom on the far right of the row of 6 is the only one to have a window. Its
square frame floods the room with natural light and allows views of the adjacent low
rise where the Musk Observatory formerly stood. Looming behind this is the side of the escarpment above, topped by a layer of sandstone that was once a seabed, as attested by the surface layer of fossilized shells.
Windows will be prized on Mars.

GreenHab Report – November 21st

Crew 215 GreenHab Report 21-11-2019
GreenHab Officer: Guy Murphy
Environmental control: Door left open for ventilation midday then reclosed 5pm.
Average temperature: 20C (temperature read 3pm)
Hours of supplemental light: Automatic system
Daily water usage for crops: 16 litres (2 full blue watering cans).
Daily water usage for research and/or other purposes:
Water in Blue Tank –
Time(s) of watering for crops: 9am, 3pm.
Changes to crops: 1 green apple seed planted, repotted aloe vera plants. Noted that 3
self seeded tomato plants have appeared in large planter tray, species unknown. (I will
transfer to pots.)

Established plants introduced 03/11/2019

[3x] strawberry plants (Everbearing), in medium metal planter
[1x] spicy orange thyme (thymus fragantissimus), small terra cotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terra cotta pot
[1x] Mint Mentha – Berries & Cream, small terra cotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terra cotta pot
[1x] Ice plant – Delospermo ‘jewel of the desert garnet’, small terra cotta pot
[1x] Sempervium textorum assorted Hen & Chicks, small terra cotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot
[2x] small Aloe Vera, in medium metal planter, frost damaged (removed 1311219)

Seeds Planted 04/11/2019 (Large metal planting tub)
Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officialnis
Sage – Savia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris

Seeds Planted 10/11/2019 (Large metal planting tub)
Corinader – Coriandrum Sativum

Seeds Planted in12/11/2019 (2 smaller planting tubs)
Carrots ‘Purple Sun F1’
Radish ‘icicle’

Seeds Planted 13/11/2019
Perpetual Spinach, Rocket, Spinach ‘Lakeside F1’in medium metal planter.
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.

Seeds Planted 14/11/2019
Tomato (Chocolate Cherry) and Tomato Pole Roma in small pots for transplantation.
Onion (Evergreen White Bunching)
Cucumbers (Patio Snackers F1) in hanging pot
Pepper (Mini Bell Color Mix) in small pots for transplantation.

Seeds Planted 15/11/2019
Acorn Pumpkin in a small pot for future transplantation. Seed taken from a pumpkin

brought to the Hab at start of rotation as a vegetable.
Snow Peas in 2 hanging pots.

Bulbs planted 19/11/2019
25 white Narcissus bulbs (Paperwhites ‘Ziva’)
Planted into 3 small terra cotta pots & 2 long planter boxes.

Seed planted 21/11/2019
1 apple seed (green apple) from a fruit brought to the campus early in the mission.

Other seeds stored in the Greenhab are identified as –

Rocket ‘Arugula Rocket Salad ‘Rocky’ (Diplotaxis tenuifolia), 1/2 packet
Spinach ‘Lakeside F1’, 1/2 packet
Carrots ‘Purple Sun F1’ – 1/2 packet
Radish ‘icicle’ – 1/2 packet
Snow Pea (Oregon Giant)
Cucumbers (Patio Snackers F1), 1/2 packet
Onion (Evergreen White Bunching), 1/2 packet
Pepper (Mini Bell Color Mix), 1/2 packet
Perpetual Spinach (Green Leaf Chard), 1/2 packet

Partly used packed of seed have been stored in the provided tin to protect from vermin.
This will also allow repropagation later in the season to stagger harvest times and
allow resowing if crops are lost.

Harvest: None

Support/supplies needed: Bag of rubbish requires removal. Would be desirable for more
wire hanging points be installed on roof frame for white IKEA hanging baskets.

Operations Report – November 21st

Crew 215 Operations Report 21-Nov-2019
SOL: 11
Name of person filing report: Shane Usher
Non-nominal systems: N/A
Notes on non-nominal systems: N/A
Generator: run
Hours run: 12 h 10 min
From what time last night: 1850
To what time this morning: 0700
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 69%
Diesel Reading – 45%
Station Propane Reading – 65%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1459976
Water (static tank) (gal): 276
Static to Loft Pump used – yes
Water in Green Hab (gal): 234.5
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: Assigned to Director
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 112.7
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes
Opportunity rover used: no
Hours: 69
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: no
Hours: 119.1
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: N/A
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used (gal): 0
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? With Director
CrewCar used and why, where? At Grand Junction awaiting Crew 216
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: N/A
Summary of Hab operations: N/A
Summary of GreenHab operations: Atmospheric equalization to the green hab carried out
during the afternoon, plants were watered (16L,4.6 gal). Door closed at 5pm, heater on
Summary of Science Dome operations: Gypsum tile synthesis experiments concluded, splint
manufacture conducted.
Summary of RAM operations: N/A
Summary of any observatory issues: Nominal
Summary of health and safety issues: Checklist controlled safety medical drill carried
Questions, concerns and requests to Mission Support: N/A

Science Report – November 21st

Science Report 21 November 2019
SOL: 11
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The experimenting with gypsum and sand to create a form of plaster of paris cement was completed. After drying, the tiles will be ready for strength testing. Approximately 10ml of water was collected as a product of the heating and will be tested as suitable for medical solutions. In the afternoon, fine grained, pure gypsum was heated and prepared as a plaster of paris and a small splint was successfully molded and used to immobilize an index finger.
Green Hab: Green Hab operations have continued with additional seeds and seedlings being planted. See Green Hab report.
RAM Operations: N/A
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin(TM) monitoring vests. A final checklist-supported response simulated emergency was carried out in the afternoon to further test the stress on untrained personnel involved in medical emergencies in isolated situations.
Additional activities: N/A

Sol Summary Report – November 21th

Crew 215 Sol Summary Report 21-NOV-2019
Sol: 11
Summary Title: Finishing up Experiments
Author’s name: Andrew Wheeler
Mission Status: Active
Sol Activity Summary: More overnight rain continued to leave the environs impassable but a strong wind in the early morning had things drying out. After the morning brief, more seedlings and seeds were planted in the Green Hab and, in the Science Dome, the making of gypsum tiles continued. The final tiles were completed in the afternoon and have been left out to dry. Just after lunch, the final emergency simulation was carried when a checklist-supported response was initiated for an injured crew member within the hab. It was successfully carried out and evaluated. Meanwhile, another test of the gypsum powder was devised when a small splint was manufactured and applied to a simulated broken index finger. Though it was applied before being fully set, the gypsum splint proved successful in immobilizing the finger for 2 hours before it was removed. A further 1.5ml of water was collected for testing as a basis for medical grade solutions. Dinner tonight was tortilla soup with fresh bread followed by chocolate pudding with jelly, fruit salad and peanut butter cookies.
Look Ahead Plan: Cleaning of staterooms and living areas, inventorying food and equipment and cleaning of the Science, Green Hab and RAM areas in preparation for end of rotation handover. Packing away of experiments and personal equipment.
Anomalies in work: N/A
Weather: Rain and clouds clearing, light winds.
Crew Physical Status: Healthy
Reports to be file: Sol Summary Report, Operations Report, Science Report, Green Hab Report, Journalist Report, Daily Photos.
Support Requested: N/A

Copyright © The Mars Society. All rights reserved. | Main Site